# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### 06 Mar 2026 v4

@@ 1.1 Backgrounds and motivation
Boson Sampling is a computational task that is complexity-theoretically proven to be hard to classically simulate under plausible assumptions [aaronson2011computational, hamilton2017gaussian, deshpande2022quantum]. Accordingly, Boson Sampling has gathered significant attention, as it would possibly play a key role in the experimental demonstration of quantum computational advantage using near-term quantum devices. However, the implementation of Boson Sampling in experimental settings with near-term quantum devices is inevitably subject to various sources of noise [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian]. The problem is that those noises would possibly rule out the classical intractability of Boson Sampling, and thus potentially hinder the experimental demonstration of quantum advantage with Boson Sampling. Indeed, both for finite-size near-term experiments and asymptotic limits as system size scales, numerous studies [renema2018classical, renema2018efficient, moylett2019classically, renema2020simulability, shi2022effect, van2024efficient, shchesnovich2019noise, oszmaniec2018classical, garcia2019simulating, qi2020regimes, brod2020classical, villalonga2021efficient, oh2021classical2, bulmer2022boundary, liu2023simulating, oh2023classical, oh2024classicalalgorithm, oh2025classical, oh2025recent] have proposed efficient classical simulation algorithms of Boson Sampling under various noise models, such as photon loss, partial distinguishability, random Gaussian noise, etc. Their results indicate that as the noise rate of Boson Sampling increases, it eventually renders such a noisy sampler classically simulable. Moreover, as the noise is typically accumulated with each circuit depth, the quantum signal for classical intractability exhibits exponential decay with increasing circuit depth. Hence, circuits with polynomially increasing depth with system size would suffer from significantly enlarged noise rates, posing substantial challenges to achieving quantum advantage in such settings.
@@ 1.2 Our results: Average-case hardness of shallow-depth Boson Sampling
Most of the current theoretical foundations of the average-case hardness of Boson Sampling rely on global Haar random unitary circuits [aaronson2011computational, hamilton2017gaussian, deshpande2022quantum, grier2022complexity, bouland2022noise, bouland2023complexity, bouland2024average], as they almost satisfy the two conditions described above. Namely, the outcome instances can be effectively hidden by global Haar random unitaries, and approximating the output probability within sufficiently large $\epsilon$ on average over global Haar random circuit instances is #P-hard under some conjectures. However, the problem for our case is that implementing a global Haar random unitary requires at least polynomially large circuit depth (e.g., see [zyczkowski1994random, russell2017direct]), and thus is not implementable in sub-polynomial circuit depths. This poses a challenge in proving the average-case hardness of Boson Sampling in the shallow-depth regime—specifically, the lack of the hiding property. Due to the absence of hiding in the shallow-depth regime, the random outcome instances of Boson Sampling cannot be hidden by random circuit instances. Accordingly, even if we establish average-case hardness over randomly chosen circuits for a fixed outcome, it does not directly imply classical simulation hardness via Stockmeyer’s reduction, which requires average-case hardness over randomly chosen outcomes [aaronson2011computational].
@@
Another problem is that there already exist efficient classical algorithms that can approximately simulate shallow-depth Boson Sampling in certain circumstances, which directly rule out the classical simulation hardness in the shallow-depth regime for those cases. While exact simulation of Boson Sampling is classically hard even for constant-depth circuits [brod2015complexity], approximate simulation becomes easy for 1-dimensional local log-depth circuits [vidal2003efficient, qi2022efficient] and also for more general dimension local circuits under some constraints [deshpande2018dynamical, oh2022classical, kolarovszki2023simulating]. According to their results, if we use circuits composed of only geometrically local gates, a polynomial circuit depth is generally required for a sufficiently large correlation to achieve approximate simulation hardness. Those results indicate that we cannot expect the hardness results in the most general case of shallow-depth circuits composed of local gates only.
@@
To deal with those problems, we take the following approach: First, we consider shallow linear optical circuit architectures that fully employ geometrically non-local gates. In fact, the implementation of non-local gates is promising for near-term experimental settings; for example, experiments of linear optical systems based on trapped ions [shen2014scalable, chen2023scalable] and photonic architecture [madsen2022quantum] implemented long-range interactions. Because of the absence of hiding, we derive the average-case hardness over both randomly chosen outcomes and randomly chosen circuits over the shallow circuit architectures, by establishing a worst-to-average-case reduction for both outcome and circuit instances. In short, this reduction process can be done in two steps: $(i)$ from a given fixed outcome to a randomly chosen collision-free outcome, and $(ii)$ from a worst-case circuit to a randomly chosen circuit over a local random circuit distribution.
@@
For random circuit ensembles, we employ a local random circuit ensemble in the shallow circuit architecture, inspired by the hardness results of random circuit sampling [bouland2019complexity, movassagh2023hardness, bouland2022noise, kondo2022quantum, krovi2022average]. Here, the local random distribution in this context means that each gate composing the circuit is independently chosen Haar random gate; we note that recent experimental setups of Boson Sampling [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian] follow a similar circuit distribution, but with geometrically local architectures. Additionally, due to the absence of symmetry, the local random circuit ensemble itself does not guarantee the bosonic birthday paradox, which is crucial for the hardness of Boson Sampling when using only collision-free outcomes [aaronson2011computational]. Hence, we also include a random permutation circuit at the input of our random circuit to guarantee the bosonic birthday paradox for our random circuit ensemble (see Appendix B). In fact, this random permutation circuit can be implemented within the shallow depth regime (see, e.g., Lemma 1), and can be experimentally realized by preparing a random input configuration instead.
@@
Also, since our average-case hardness result regards both the random outcomes and random circuits, it is not straightforward to show the classical simulation hardness of shallow-depth Boson Sampling as in the original Boson Sampling proposal [aaronson2011computational]. Accordingly, we show how our average-case hardness result over both the randomly chosen outcomes and circuits leads to the classical simulation hardness argument. This implies that improving the additive imprecision for our average-case hardness result is the only remaining problem for the fully theoretically guaranteed classical intractability of shallow-depth Boson Sampling.
@@ 2.1 Notations
Let us define the total mode number as $M$, where we set $M$ as a power of 2 for simplicity. We set the output photon number $N$ polynomially related to $M$ as $M=c_{0}N^{\gamma}$ for a constant $c_{0}$ and $\gamma\geq 2$, to ensure that collision-free outcomes dominate the probability weight over all possible $N$-photon outcomes over $M$ modes (i.e., by bosonic birthday paradox; see Refs. [aaronson2011computational, arkhipov2012bosonic] and Appendix B). We use the notation $\bm{s}$ as an $M$-dimensional output configuration vector representing $N$-photon outcome over $M$ modes, such that each element $s_{i}$ ($i\in[M]$) of $\bm{s}$ denotes the occupied number of photons in the $i$th mode. Namely, $\bm{s}$ can be expressed as $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in[N]$ with the constraint $\sum_{i=1}^{M}s_{i}=N$. Also, let $\bm{t}$ be an $M$-dimensional input configuration vector corresponding to an $N$-photon input state over $M$ modes, such that $\bm{t}=(t_{1},\dots,t_{M})$ where now each $t_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}t_{i}=N$.
Hereafter, we use the terminology Boson Sampling representing Boson Sampling with Fock-state inputs as in the original Boson Sampling proposal [aaronson2011computational]; we will explicitly use the terminology Gaussian Boson Sampling for Boson Sampling with Gaussian state inputs [hamilton2017gaussian]. We now define $p_{\bm{s}}(C)$ as the output probability of Boson Sampling with a linear optical circuit (unitary) matrix $C$ to obtain an outcome $\bm{s}$ from a fixed input configuration $\bm{t}$. More precisely, $p_{\bm{s}}(C)$ can be represented as [aaronson2011computational]
@@
In this work, we consider the shallow-depth circuit architecture, specifically in logarithmic depth regime $D=O(\log N)$. As we only consider the binary logarithm throughout this work, we will omit the base of the logarithm such that $\log(\cdot)$ implies $\log_{2}(\cdot)$ hereafter. We now define the shallow linear optical circuit architecture of circuit depth $D=\log M$ with the number of gates $m=\frac{M}{2}\log M$, using the convention used in [dao2020kaleidoscope, go2024exploring].
@@
Next, we define the Kaleidoscope circuit architecture proposed in [dao2020kaleidoscope], using the butterfly circuit architecture defined above.
@@ 2.2 Definitions for random circuit and outcome ensemble
Indeed, this circuit ensemble is similar to the random circuit ensemble employed by the previous random circuit sampling proposals [bouland2019complexity, movassagh2023hardness, bouland2022noise, kondo2022quantum, krovi2022average] which employ gate-wise Haar random unitary circuits. Next, we introduce our second random circuit ensemble, which is the random permutation circuit ensemble over the fixed shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
@@
Note that this random permutation circuit ensemble can be implemented in the shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$ by Lemma 1. Before proceeding, we briefly remark on the motivation for including the permutation circuit ensemble in our random circuit construction. Specifically, for any $M$-mode circuit architecture $\mathcal{A}$, including $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ considered in this work, a circuit drawn from the product circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ guarantees the bosonic birthday paradox [aaronson2011computational, arkhipov2012bosonic]. The bosonic birthday paradox ensures that the probability of obtaining only collision outcomes is sufficiently large given that the photon number $N$ is sufficiently smaller than the mode number $M$, as stated in the following lemma.
@@ 3 Worst-case hardness of output probability estimation
We briefly sketch the proof of our worst-case hardness result for the shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$. The proof is based on the result by [brod2015complexity], which showed the simulation hardness of exact Boson Sampling with constant-depth linear optical circuits. Specifically, as shown in [brod2015complexity], based on measurement-based quantum computation (MBQC) scheme, there exists a constant-depth linear optical circuit that can simulate an arbitrary given quantum circuit using post-selection. To extend this result to our case, we show that the constant-depth worst-case MBQC circuit in [brod2015complexity] can be embedded in our circuit architecture $\mathcal{B}\mathcal{B}^{*}$. Combining these arguments, additively approximating the output probability of any quantum circuit can be reduced to additively approximating the output probability of any circuit in $\mathcal{B}\mathcal{B}^{*}$, with imprecision blowup up to the inverse of post-selection probability. Using the fact that the additive approximation of any quantum circuit is #P-hard for certain additive imprecision [kondo2022quantum], we can obtain the worst-case hardness of our shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
@@
We conclude this section by noting that, it is not necessary to derive worst-case hardness of shallow-depth Boson Sampling exclusively for the circuit architecture $\mathcal{BB}^{*}$. That is, embedding the constant-depth worst-case MBQC circuit [brod2015complexity] into the logarithmic-depth $\mathcal{BB}^{*}$ circuit, as done in our proof of Theorem 3, is not essential for the worst-case hardness of shallow-depth Boson Sampling in principle (and, as will be discussed later, not even for the worst-to-average-case reduction). Although not mandatory, however, since we later make use of $\mathcal{BB}^{*}$ to implement the global permutation during the worst-to-average-case reduction (due to the absence of hiding), we establish worst-case hardness within $\mathcal{BB}^{*}$ so as to simplify the overall circuit architecture to the repeated structure $(\mathcal{BB}^{*})^{q}$ defined in Definition 3.
@@ 4 Average-case hardness of output probability estimation
First, since our average-case argument regards both outcomes and circuits, we begin by describing how to fix the outcome in establishing the worst-to-average-case reduction. The remaining problem is then to establish a “circuit-level" worst-to-average-case reduction, from a worst-case circuit to randomly chosen average-case circuits for a fixed output probability, similarly to previous hardness results of random circuit sampling [bouland2019complexity, movassagh2023hardness, bouland2022noise, kondo2022quantum, krovi2022average]. To fix the outcome, our strategy is to randomly permute both a given worst-case circuit $C_{0}$ in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$ and a given collision-free outcome $s_{0}$. That is, we sample random $M$-mode permutation $\bm{P}_{0}$ uniformly, and permute the worst-case circuit $C_{0}$ and the fixed outcome $\bm{s}_{0}$ in Theorem 3 equally with $\bm{P}_{0}$. Then, the permuted outcome $\bm{s}=\bm{P}_{0}\bm{s}_{0}$ essentially follows the random outcome ensemble $\bm{s}\sim\mathcal{G}_{M,N}$, and the permuted circuit $\bm{P}_{0}C_{0}$ is in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ by Lemma 1. Note that the fixed worst-case output probability $p_{\bm{s}_{0}}(C_{0})$ is now equivalent to $p_{\bm{s}}(\bm{P}_{0}C_{0})$, and thus our new goal is to estimate $p_{\bm{s}}(\bm{P}_{0}C_{0})$ value, given estimated average-case values of $p_{\bm{s}}(U)$.
@@
We also remark again that one could further reduce the circuit depth by taking the worst-case circuit $C_{0}$ as the constant-depth MBQC circuit in [brod2015complexity] directly, without encoding into the logarithmic-depth architecture $(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$ as in Theorem 3. In this case, the overall circuit architecture $\mathcal{A}$ for the average-case hardness consists of the constant-depth circuit (for MBQC) sandwiched between the $\mathcal{BB}^{*}$ architectures used for implementing required permutations (i.e., $\bm{P}_{0}$ and $\bm{P}_{1}$), yielding a slightly shallower circuit architecture than $\mathcal{A}=(\mathcal{BB}^{*})^{q\geq 3}$ used in our case. Nevertheless, since we are already in the logarithmic-depth regime (by implementing permutations with $\mathcal{BB}^{*}$), we adopt $\mathcal{A}=(\mathcal{BB}^{*})^{q\geq 3}$ as our overall circuit architecture for the ensemble $\mathcal{H}_{\mathcal{A}}$ in order to provide a unified description of the circuit architecture required for the hardness of shallow-depth Boson Sampling, as well as a clearer layout for experimental implementations.

### 11 Mar 2026 v5

@@ 1.1 Backgrounds and motivation
Boson Sampling is a computational task that is complexity-theoretically proven to be hard to classically simulate under plausible assumptions [1, 2, 3]. Accordingly, Boson Sampling has gathered significant attention, as it would possibly play a key role in the experimental demonstration of quantum computational advantage using near-term quantum devices. However, the implementation of Boson Sampling in experimental settings with near-term quantum devices is inevitably subject to various sources of noise [4, 5, 6, 7]. The problem is that those noises would possibly rule out the classical intractability of Boson Sampling, and thus potentially hinder the experimental demonstration of quantum advantage with Boson Sampling. Indeed, both for finite-size near-term experiments and asymptotic limits as system size scales, numerous studies [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] have proposed efficient classical simulation algorithms of Boson Sampling under various noise models, such as photon loss, partial distinguishability, random Gaussian noise, etc. Their results indicate that as the noise rate of Boson Sampling increases, it eventually renders such a noisy sampler classically simulable. Moreover, as the noise is typically accumulated with each circuit depth, the quantum signal for classical intractability exhibits exponential decay with increasing circuit depth. Hence, circuits with polynomially increasing depth with system size would suffer from significantly enlarged noise rates, posing substantial challenges to achieving quantum advantage in such settings.
@@ 1.2 Our results: Average-case hardness of shallow-depth Boson Sampling
Most of the current theoretical foundations of the average-case hardness of Boson Sampling rely on global Haar random unitary circuits [1, 2, 3, 27, 28, 29, 30], as they almost satisfy the two conditions described above. Namely, the outcome instances can be effectively hidden by global Haar random unitaries, and approximating the output probability within sufficiently large $\epsilon$ on average over global Haar random circuit instances is #P-hard under some conjectures. However, the problem for our case is that implementing a global Haar random unitary requires at least polynomially large circuit depth (e.g., see [31, 32]), and thus is not implementable in sub-polynomial circuit depths. This poses a challenge in proving the average-case hardness of Boson Sampling in the shallow-depth regime—specifically, the lack of the hiding property. Due to the absence of hiding in the shallow-depth regime, the random outcome instances of Boson Sampling cannot be hidden by random circuit instances. Accordingly, even if we establish average-case hardness over randomly chosen circuits for a fixed outcome, it does not directly imply classical simulation hardness via Stockmeyer’s reduction, which requires average-case hardness over randomly chosen outcomes [1].
@@
Another problem is that there already exist efficient classical algorithms that can approximately simulate shallow-depth Boson Sampling in certain circumstances, which directly rule out the classical simulation hardness in the shallow-depth regime for those cases. While exact simulation of Boson Sampling is classically hard even for constant-depth circuits [33], approximate simulation becomes easy for 1-dimensional local log-depth circuits [34, 35] and also for more general dimension local circuits under some constraints [36, 37, 38]. According to their results, if we use circuits composed of only geometrically local gates, a polynomial circuit depth is generally required for a sufficiently large correlation to achieve approximate simulation hardness. Those results indicate that we cannot expect the hardness results in the most general case of shallow-depth circuits composed of local gates only.
@@
To deal with those problems, we take the following approach: First, we consider shallow linear optical circuit architectures that fully employ geometrically non-local gates. In fact, the implementation of non-local gates is promising for near-term experimental settings; for example, experiments of linear optical systems based on trapped ions [39, 40] and photonic architecture [6] implemented long-range interactions. Because of the absence of hiding, we derive the average-case hardness over both randomly chosen outcomes and randomly chosen circuits over the shallow circuit architectures, by establishing a worst-to-average-case reduction for both outcome and circuit instances. In short, this reduction process can be done in two steps: $(i)$ from a given fixed outcome to a randomly chosen collision-free outcome, and $(ii)$ from a worst-case circuit to a randomly chosen circuit over a local random circuit distribution.
@@
For random circuit ensembles, we employ a local random circuit ensemble in the shallow circuit architecture, inspired by the hardness results of random circuit sampling [41, 42, 28, 43, 44]. Here, the local random distribution in this context means that each gate composing the circuit is independently chosen Haar random gate; we note that recent experimental setups of Boson Sampling [4, 5, 6, 7] follow a similar circuit distribution, but with geometrically local architectures. Additionally, due to the absence of symmetry, the local random circuit ensemble itself does not guarantee the bosonic birthday paradox, which is crucial for the hardness of Boson Sampling when using only collision-free outcomes [1]. Hence, we also include a random permutation circuit at the input of our random circuit to guarantee the bosonic birthday paradox for our random circuit ensemble (see Appendix B). In fact, this random permutation circuit can be implemented within the shallow depth regime (see, e.g., Lemma 1), and can be experimentally realized by preparing a random input configuration instead.
@@
Also, since our average-case hardness result regards both the random outcomes and random circuits, it is not straightforward to show the classical simulation hardness of shallow-depth Boson Sampling as in the original Boson Sampling proposal [1]. Accordingly, we show how our average-case hardness result over both the randomly chosen outcomes and circuits leads to the classical simulation hardness argument. This implies that improving the additive imprecision for our average-case hardness result is the only remaining problem for the fully theoretically guaranteed classical intractability of shallow-depth Boson Sampling.
@@ 2.1 Notations
Let us define the total mode number as $M$, where we set $M$ as a power of 2 for simplicity. We set the output photon number $N$ polynomially related to $M$ as $M=c_{0}N^{\gamma}$ for a constant $c_{0}$ and $\gamma\geq 2$, to ensure that collision-free outcomes dominate the probability weight over all possible $N$-photon outcomes over $M$ modes (i.e., by bosonic birthday paradox; see Refs. [1, 45] and Appendix B). We use the notation $\bm{s}$ as an $M$-dimensional output configuration vector representing $N$-photon outcome over $M$ modes, such that each element $s_{i}$ ($i\in[M]$) of $\bm{s}$ denotes the occupied number of photons in the $i$th mode. Namely, $\bm{s}$ can be expressed as $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in[N]$ with the constraint $\sum_{i=1}^{M}s_{i}=N$. Also, let $\bm{t}$ be an $M$-dimensional input configuration vector corresponding to an $N$-photon input state over $M$ modes, such that $\bm{t}=(t_{1},\dots,t_{M})$ where now each $t_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}t_{i}=N$.
Hereafter, we use the terminology Boson Sampling representing Boson Sampling with Fock-state inputs as in the original Boson Sampling proposal [1]; we will explicitly use the terminology Gaussian Boson Sampling for Boson Sampling with Gaussian state inputs [2]. We now define $p_{\bm{s}}(C)$ as the output probability of Boson Sampling with a linear optical circuit (unitary) matrix $C$ to obtain an outcome $\bm{s}$ from a fixed input configuration $\bm{t}$. More precisely, $p_{\bm{s}}(C)$ can be represented as [1]
@@
In this work, we consider the shallow-depth circuit architecture, specifically in logarithmic depth regime $D=O(\log N)$. As we only consider the binary logarithm throughout this work, we will omit the base of the logarithm such that $\log(\cdot)$ implies $\log_{2}(\cdot)$ hereafter. We now define the shallow linear optical circuit architecture of circuit depth $D=\log M$ with the number of gates $m=\frac{M}{2}\log M$, using the convention used in [46, 47].
@@
Next, we define the Kaleidoscope circuit architecture proposed in [46], using the butterfly circuit architecture defined above.
@@ 2.2 Definitions for random circuit and outcome ensemble
Indeed, this circuit ensemble is similar to the random circuit ensemble employed by the previous random circuit sampling proposals [41, 42, 28, 43, 44] which employ gate-wise Haar random unitary circuits. Next, we introduce our second random circuit ensemble, which is the random permutation circuit ensemble over the fixed shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
@@
Note that this random permutation circuit ensemble can be implemented in the shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$ by Lemma 1. Before proceeding, we briefly remark on the motivation for including the permutation circuit ensemble in our random circuit construction. Specifically, for any $M$-mode circuit architecture $\mathcal{A}$, including $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ considered in this work, a circuit drawn from the product circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ guarantees the bosonic birthday paradox [1, 45]. The bosonic birthday paradox ensures that the probability of obtaining only collision outcomes is sufficiently large given that the photon number $N$ is sufficiently smaller than the mode number $M$, as stated in the following lemma.
@@ 3 Worst-case hardness of output probability estimation
We briefly sketch the proof of our worst-case hardness result for the shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$. The proof is based on the result by [33], which showed the simulation hardness of exact Boson Sampling with constant-depth linear optical circuits. Specifically, as shown in [33], based on measurement-based quantum computation (MBQC) scheme, there exists a constant-depth linear optical circuit that can simulate an arbitrary given quantum circuit using post-selection. To extend this result to our case, we show that the constant-depth worst-case MBQC circuit in [33] can be embedded in our circuit architecture $\mathcal{B}\mathcal{B}^{*}$. Combining these arguments, additively approximating the output probability of any quantum circuit can be reduced to additively approximating the output probability of any circuit in $\mathcal{B}\mathcal{B}^{*}$, with imprecision blowup up to the inverse of post-selection probability. Using the fact that the additive approximation of any quantum circuit is #P-hard for certain additive imprecision [43], we can obtain the worst-case hardness of our shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
@@
We conclude this section by noting that, it is not necessary to derive worst-case hardness of shallow-depth Boson Sampling exclusively for the circuit architecture $\mathcal{BB}^{*}$. That is, embedding the constant-depth worst-case MBQC circuit [33] into the logarithmic-depth $\mathcal{BB}^{*}$ circuit, as done in our proof of Theorem 3, is not essential for the worst-case hardness of shallow-depth Boson Sampling in principle (and, as will be discussed later, not even for the worst-to-average-case reduction). Although not mandatory, however, since we later make use of $\mathcal{BB}^{*}$ to implement the global permutation during the worst-to-average-case reduction (due to the absence of hiding), we establish worst-case hardness within $\mathcal{BB}^{*}$ so as to simplify the overall circuit architecture to the repeated structure $(\mathcal{BB}^{*})^{q}$ defined in Definition 3.
@@ 4 Average-case hardness of output probability estimation
First, since our average-case argument regards both outcomes and circuits, we begin by describing how to fix the outcome in establishing the worst-to-average-case reduction. The remaining problem is then to establish a “circuit-level" worst-to-average-case reduction, from a worst-case circuit to randomly chosen average-case circuits for a fixed output probability, similarly to previous hardness results of random circuit sampling [41, 42, 28, 43, 44]. To fix the outcome, our strategy is to randomly permute both a given worst-case circuit $C_{0}$ in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$ and a given collision-free outcome $s_{0}$. That is, we sample random $M$-mode permutation $\bm{P}_{0}$ uniformly, and permute the worst-case circuit $C_{0}$ and the fixed outcome $\bm{s}_{0}$ in Theorem 3 equally with $\bm{P}_{0}$. Then, the permuted outcome $\bm{s}=\bm{P}_{0}\bm{s}_{0}$ essentially follows the random outcome ensemble $\bm{s}\sim\mathcal{G}_{M,N}$, and the permuted circuit $\bm{P}_{0}C_{0}$ is in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ by Lemma 1. Note that the fixed worst-case output probability $p_{\bm{s}_{0}}(C_{0})$ is now equivalent to $p_{\bm{s}}(\bm{P}_{0}C_{0})$, and thus our new goal is to estimate $p_{\bm{s}}(\bm{P}_{0}C_{0})$ value, given estimated average-case values of $p_{\bm{s}}(U)$.
@@
We also remark again that one could further reduce the circuit depth by taking the worst-case circuit $C_{0}$ as the constant-depth MBQC circuit in [33] directly, without encoding into the logarithmic-depth architecture $(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$ as in Theorem 3. In this case, the overall circuit architecture $\mathcal{A}$ for the average-case hardness consists of the constant-depth circuit (for MBQC) sandwiched between the $\mathcal{BB}^{*}$ architectures used for implementing required permutations (i.e., $\bm{P}_{0}$ and $\bm{P}_{1}$), yielding a slightly shallower circuit architecture than $\mathcal{A}=(\mathcal{BB}^{*})^{q\geq 3}$ used in our case. Nevertheless, since we are already in the logarithmic-depth regime (by implementing permutations with $\mathcal{BB}^{*}$), we adopt $\mathcal{A}=(\mathcal{BB}^{*})^{q\geq 3}$ as our overall circuit architecture for the ensemble $\mathcal{H}_{\mathcal{A}}$ in order to provide a unified description of the circuit architecture required for the hardness of shallow-depth Boson Sampling, as well as a clearer layout for experimental implementations.